OpenStudy (gabylovesyou):

Brandon found that a solution of x = 1 was extraneous for his function. What would be a possible denominator in Brandon's function? 2x + 4 x2 + 3x + 9 3x − 11 x2 + 3x − 4

OpenStudy (gabylovesyou):

@welshfella

OpenStudy (kendricklamar2014):

x=1 must be a root of the denominator. Since the denominator cannot be zero, it is undefined at this point. Answer: x2+3x-4=1+3-4=0 is the equation that youre looking for

OpenStudy (gabylovesyou):

thank you

OpenStudy (kendricklamar2014):

Your welcome